Megan factored the expression -12x^2+52x-35 as (-2x+5)(6x-7) But when Jacob applied the FOIL principle and multiplied out (-2x+5)(6x-7) he got -12x^2+44x-35; thus, Megan’s solution does not appear to check. Why is that? Please help Megan to understand this better. Explain your reasoning and correctly factor the original expression, if possible. If the expression is prime, so state.
Answer by lenny460(1073)

Factor:

-12x^2 + 52x - 35

Factor a trinomial:

-(6x - 5)(2x -7)

Multiply the above using FOIL Method:
6x * 2x = 12x^2
6x * -7 = -42x
-5 * 2x = -10x
-5 * -7 = 35

12x^2 - 42x - 10x + 35
Combine like terms:
-42x - 10x = -52x

-1(12x^2 - 52x + 35)
-1 * 12x^2 = -12x^2
-1 * -52x = 52x
-1 * 35 = -35

We therefore get:

-12x^2 + 52x - 35

The Factors are:

-(6x - 5)(2x - 7)


There is another way to factor:

-12x^2 + 52x - 35
The factors are:

(-7 + 2x)(5 - 6x)

Let's check by the FOIL Method:

-7 * 5 = -35
-7 * -6x = 42x
2x * 5 = 10x
2x * -6x = -12x^2

-35 + 42x + 10 x - 12x^2

Combine like terms:
42x + 10x = 52x

-35 + 52x - 12x^2 or

-12x^2 + 52x - 35


obuong3@aol.com